# Solving High-dimensional Inverse Problems Using Amortized Likelihood-free Inference with Noisy and Incomplete Data

### v1

@@ Abstract
We present a novel likelihood-free probabilistic inversion method based on normalizing flows for high-dimensional inverse problems. The proposed method is comprised of two complementary networks: a summary network for data compression, and an inference network for parameter estimation. The summary network encodes raw observations into a fixed-size vector of summary statistics, while the inference network generates samples of the approximate posterior distribution of the model parameters based on these summary statistics. The posterior samples are produced in a deep generative fashion by sampling from a latent Gaussian distribution and passing these samples through an invertible transformation. We construct this invertible transformation by sequentially alternating conditional invertible neural network (cINN) and conditional neural spline flow (cNSF) layers. The summary and inference networks are trained simultaneously.
We apply the proposed method to an inversion problem in groundwater hydrology to estimate the posterior distribution of the system’s log-conductivity field conditioned on spatially sparse time-series observations of the system’s hydraulic head responses. The conductivity field is represented with 706 degrees of freedom in the considered problem. The comparison with the likelihood-based iterative ensemble smoother PEST-IES method demonstrates that the proposed method accurately estimates the parameter posterior distribution and the observations’ predictive posterior distribution at a fraction of the inference time of PEST-IES.
@@ 1 Introduction
Inverse problems are key for characterizing parameterized models of physical systems. Unlike forward problems, which predict outcomes given model parameters, inverse problems work in the reverse direction, seeking to estimate model parameters or system states from partial measurements of observable quantities [1]. Inverse problems allow us to update computational models to better align with observed data, thereby enhancing the models’ predictive capabilities [2]. Application areas leveraging inverse problems include medical imaging [3], geophysics [4], structural health monitoring [5], reservoir engineering [6], astronomy [7], among others.
@@
BI methods can be categorized based on their approach to evaluating the likelihood function linking model parameters to observations: (i) likelihood-based inference and (ii) likelihood-free inference—also known as simulation-based inference. Likelihood-based inference estimates the posterior distribution by explicitly evaluating the likelihood function, which should be explicitly defined and tractable. The gold standard for numerically sampling the posterior distribution is Markov chain Monte Carlo (MCMC) method. While conceptually straightforward, likelihood-based inference poses various challenges. First, often directly evaluating the likelihood function is either intractable or unfeasible [12] as in the case of multi-level models where sub-models are hierarchically connected [13], continuous-time state-space models [14], and nonlinear dynamic models [15], among others. Second, likelihood-based inference requires the repeated evaluation of the forward model to ensure convergence to the true posterior, which can be computationally intractable depending on the computational cost of the forward model. Third, likelihood-based inference methods are often hampered by the curse of dimensionality, which is potentially challenging for problems with spatially heterogeneous parameters, which often demand high-dimensional characterizations [16]. While the use of surrogate models, such as Gaussian processes [17, 18], Polynomial Chaos Expansion [19, 20], and Deep Neural Network [21, 22] models, can reduce the computational cost, these methods still require likelihood evaluation and introduce additional epistemic uncertainty. Moreover, constructing accurate surrogate models becomes particularly problematic when dealing with high-dimensional inputs and outputs.
@@
In contrast, likelihood-free inference aims to estimate the posterior distribution without directly evaluating the likelihood function and relying solely on synthetic data generated for various parameter choices. Approximate Bayesian Computation (ABC) is perhaps the most developed form of likelihood-free inference [23]. In ABC, parameters are repeatedly sampled from a prior distribution, and synthetic data is generated by running the forward model using these sampled parameters. Parameters whose simulated data sufficiently resemble actual observations, as decided by user-defined criteria, are retained as samples of the target posterior; others are discarded. However, ABC has several disadvantages resulting from its accept-reject mechanism: it is challenging to set reasonable criteria for accurate posterior estimation; additionally, the entire ABC process must be restarted with new measurement data sets, limiting its application to inference for a single set of observed data [24].
@@
In recent developments, deep generative models (DGMs) have emerged as a promising approach for solving inverse problems via likelihood-free inference. Central to DGMs is the strategy of a non-linear map from a latent distribution, such as a Gaussian, to a complex target distribution. Consequently, approximating the posterior can be done by learning this non-linear transformation. Such a learning process can be efficiently executed using neural networks trained via back propagation [25]. After sufficient training, one can draw samples from a straightforward latent distribution and produce posterior distribution samples by applying the learned transformation. Numerous studies have explored the DGMs use for inverse problems, including applications of generative adversarial networks (GANs) [26, 27] and variational autoencoders [28, 29]. GANs and their variants, in particular, have become increasingly popular for likelihood-free inference. Baptista et al. [30] introduced conditional sampling for likelihood-free inference using a monotone GAN, which learns a block triangle transport map between a reference and target measure through adversarial learning combined with a monotonicity penalty. Ramesh et al. [31] developed a conditional GAN to redefine the variational objective for learning an implicit posterior distribution over latent variables, demonstrating significant promise for high-dimensional inverse problems. Additionally, Patel et al. [32] proposed using a GAN to model a field characterized by numerous discrete representations, where the GAN’s generator maps a latent distribution to the target distribution for the field. Despite the significant impact of GANs on posterior estimation, they are notoriously challenging to train, highly sensitive to hyperparameter settings, and prone to the issue of mode collapse [33].
@@
An alternative approach to DGM-based likelihood-free inference is the use of flow-based models, particularly normalizing flows [34]. Normalizing flow use invertable networks allowing direct sampling from complex, parameterized probability distributions and precise density evaluation using a change-of-variables formula. Normalizing flows have demonstrated strong performance in high-quality posterior estimation across various fields such as economics [35], epidemiology [36], physics [37], computer vision [38], and civil engineering [39]. Additionally, they provide precise log-likelihood evaluations and avoid the training stability issues often encountered in GANs.
@@
In this study, we introduce a likelihood-free inference approach using normalizing flows for estimating high-dimensional spatially heterogeneous parameters in initial-boundary value problems (IBVPs). The proposed model consists of two complementary deep neural networks: a summary network and an inference network. The summary network compresses raw time-series observation data, automatically extracting the most informative statistics and moving beyond manually crafted summary statistics. The inference network is constructed by alternating between two types of flow-based invertible layers: affine coupling layers (ACLs) and spline layers. Specifically, it integrates two types of invertible transformations: a conditional invertible neural network (cINN), which consists of a series of affine coupling layers, and a neural spline flow (cNSF), which employs spline-based transformations, such as piecewise polynomials. By alternating these layers, the inference network capitalizes on the strengths of both the ACL-based cINN and the spline-based cNSF, enhancing the model’s expressiveness and flexibility. The summary network forwards the summarized observations to the inference network, thereby conditioning the entire normalizing flow on these observations. The inference and summary networks are jointly trained through backpropagation using synthetic data. This concurrent training ensures that the data representation is adequate for estimating the posterior distributions of the parameters. We apply our approach to estimating a hydraulic log-conductivity field in a numerical groundwater model with 706 special computational cells using noisy time-dependent observations of the hydraulic head at sparse 13 locations. The estimation results are compared with the likelihood-based iterative ensemble smoother IES method implemented in the PEST software in terms of both accuracy and computational efficiency. The main advantages of the proposed method are:
@@
    it can accurately estimate high-dimensional parameters without the need to evaluate the likelihood function and make any assumptions about the posterior,
@@
  3. 3.
    it can manage datasets of varying sizes and lengths without requiring retraining.
@@ 2.1 Inverse problem formulation
The motivating application is estimation of spatially heterogeneous parameters in groundwater flow models. In this work we consider two-dimensional groundwater flow over the simulation domain $\Omega\in\mathbb{R}^{2}$ governed by the non-linear partial differential equations
$S_{y}u(\mathbf{x},t)\frac{\partial u(\mathbf{x},t)}{\partial t}=\nabla\cdot[\kappa(\mathbf{x})u(\mathbf{x},t)\nabla u(\mathbf{x},t)]-u(\mathbf{x},t)r(\mathbf{x},t),\quad\mathbf{x}\in\Omega,\,t>0$ (1)
subject to appropriate initial and boundary conditions. Here, $u(\mathbf{x},t)$ denotes the hydraulic head field and $\kappa(\mathbf{x})$ is the spatially varying hydraulic conductivity, $r(\mathbf{x},t)$ is the source term, and $S_{y}$ is the (assumed to be known) specific yield.
For a given $\kappa(\mathbf{x})$, Eq (1) is solved using the control volume finite-difference solver MODFLOW [40]. In the numerical model the domain is discretized with $N$ cells and the conductivity field is represented with the vector $\mathbf{\kappa}\in\mathbb{R}^{N}$ of its values at the cell centers. We solve the inverse problem for $\mathbf{y}=\ln\bm{\kappa}\in\mathbb{R}^{N}$ rather than $\mathbf{\kappa}$, which guarantees that $\kappa(\mathbf{x})$ is strictly positive.
We assume that only hydraulic head responses $\mathbf{u}$, collected at $N_{u}$ sensor locations ($N_{u}<N$), are available for solving the inverse problem. Since the number of observations is smaller than the number of parameters, this inverse problem is, in general, ill-posed. We note that many inverse methods reduce the dimensionality of the parameter space because of the high computational cost of finding the inverse solution in the full parameter space, an approximation leading to some errors. Here, we formulate the inverse problem in the original $\mathbb{R}^{N}$ space, which makes it possible to solve problems where the dimension reduction is not possible, i.e., when the all parameters are uncorrelated and equally important.
The proposed likelihood-free inference model is trained using a synthetic labeled dataset consisting of pairs of hydraulic head values at the sensor locations and $\mathbf{y}$. The $\mathbf{y}$ vectors are generated by sampling the log-conductivity prior distribution and the corresponding hydraulic head field is computed with MODFLOW.
@@ 2.2 Bayesian inference
Let $\bm{\theta}\coloneqq\mathbf{y}$ denote the uncertain model parameters. Furthermore, let $\mathcal{M}_{k}(\bm{\theta})$ denote the time-series forward model’s output at timestep $k$ consisting of the hydraulic head at $N_{u}$ observation locations. We also consider the excitations, e.g., boundary conditions, initial conditions, body forces, as fixed in this setting. We model the time-series observations of hydraulic head at sensor locations via the additive noise model
@@
The posterior distribution $p(\bm{\theta}|\mathbf{u}_{o,1:k})$ can be estimated through various Bayesian inference techniques, including MCMC sampling methods and ABC. However, these methods often face challenges such as high computational cost and inefficiency in high-dimensional parameter spaces, as discussed in Sec. 1. We propose a likelihood-free inference approach to address these limitations. Unlike most existing likelihood-free inference approaches, which are typically limited to low-dimensional parameter spaces (e.g., a few or up to a dozen parameter dimensions), our approach leverages the power of normalizing flows to infer high-dimensional parameter fields, specifically a 706-dimensional log-conductivity field, using sparse and noisy time-series observations. Furthermore, we introduce the integration of two complementary flow-based neural networks—cINN and cNSF—to enhance estimation accuracy. This integration allows us to alternate between layers in cINN and cNSF,enabling a more expressive and flexible mapping of the posterior distribution. The proposed framework effectively tackles the challenges of estimating high-dimensional, spatially distributed parameter fields, offering a more scalable and practical solution as detailed in the following section.
@@ 3 Methodology
The proposed likelihood-free method for parameter estimation consists of three distinct phases, as depicted in Fig. 1. In Phase I, simulation data are generated by sampling model parameters from the prior distribution, $\bm{\theta}\sim p(\bm{\theta})$, and using them to run a forward model. This process produces a training dataset of $M$ pairs, $\{(\bm{\theta}^{(i)},\mathbf{u}^{(i)}_{\text{syn},1:k})\}_{i=1}^{M}$, where $\bm{\theta}^{(i)}$ represents the sampled parameters and $\mathbf{u}^{(i)}_{\text{syn},1:k}$ denotes the corresponding simulated observations over the time steps $1:k$. Phase II consists of the offline training of the normalizing flow model using simulation data. Finally, Phase III consists of online amortized inference, in which we generate samples from the posterior distribution for any observation dataset. The term “online amortized inference” refers to the fact that in our methodology the intensive computational effort of data simulation and model training are front-loaded during phases I and II. Consequently, inference can be performed by sampling from latent simple distribution and learnt transformation, without starting anew for different sets of observation data, allowing for near-instantaneous results, often within seconds. This confers a substantial boost in efficiency, especially for real-world applications.
It is worth mentioning that the need for retraining flow model depends on the nature of the changes to $N_{u}$ (number of observation locations) and $k$ (number of time steps): If the new observations contain fewer time steps than the training data, the model does not necessarily need to be retrained. It can still be used to infer parameters by leveraging the available data. Similarly, if the number of time steps in the new observations exceeds those in the training data, the model can also be reused by employing recursive likelihood-free inference [13] to incorporate additional temporal information without requiring retraining. If the number of observation locations differs from that in the training data, retraining the model is necessary. The spatial arrangement of observations directly impacts the structure of the likelihood-free inference process and the mappings learned by the normalizing flow model.
Figure 1: The workflow of the proposed method for Bayesian model inversion
@@ 3.1 Likelihood-free inference by normalizing flow
The proposed likelihood-free inference framework is based on Normalizing Flows (NF) and integrates concepts from Bayesian inference and NF theory. Originally developed by Tabak and Turner [41], and further by Tabak and Vanden-Eijnden [42], NF was primarily applied to feature extraction and density estimation. NF is a generative model that constructs complex distributions by transforming a known simple probability distribution (e.g., a multivariate normal distribution) through a series of invertible and differentiable transformations. A notable strength of NF is its scalability, enabling it to efficiently manage distributions from simple, low-dimensional to complex, high-dimensional forms.
@@
NF consists of a sequence of $t$ transformations, applied to latent variables $\mathbf{z}$, as follows
@@
$\bm{\theta}=\mathbf{g}_{1}\circ\mathbf{g}_{2}\circ\dots\circ\mathbf{g}_{t}(\mathbf{z}_{0})$ (7)
Thus, the distribution $p(\bm{\theta})$ can be expressed in terms of the base distribution $p(\mathbf{z}_{0})$ and the Jacobian determinant of the composite transformation $\mathbf{h}(\bm{\theta})$. The complex target distribution is thus approximated by the transformed latent distribution $p(\mathbf{z}_{0})$ using change-of-variable rule. The distribution $p(\bm{\theta})$ is given by
$p(\bm{\theta})=p(\mathbf{z}_{0})\left|\det\left(\frac{\partial\mathbf{h}(\bm{\theta})}{\partial\bm{\theta}}\right)\right|=p(\mathbf{z}_{0})\left|\det\left(\frac{\partial\mathbf{g}(\mathbf{h}(\bm{\theta}))}{\partial\mathbf{z}_{0}}\right)\right|^{-1}$ (8)
@@
Figure 2: Schematic of normalizing flow
@@
$\bm{\theta}\sim p_{\bm{\alpha}}(\bm{\theta}|\mathbf{u}_{o,1:k})\Leftrightarrow\bm{\theta}=\mathbf{h}_{\bm{\alpha}}^{-1}(\mathbf{z};\mathbf{u}_{o,1:k}),\text{ with }\mathbf{z}\sim\mathcal{N}(\mathbf{z}|\bm{0},\mathbf{I})$ (12)
In practice, the design of invertible functions $\mathbf{g}(\cdot)$ for nonlinear bijective mappings poses significant challenges. Commonly, these functions are parameterized through layers of invertible neural networks, each providing efficient bijections with cost-effective Jacobian evaluations. A detailed review of such architectures is available in [43]. In this research, we leverage the conditional invertible neural network (cINN), characterized by the affine coupling layers, and the conditional neural spline flow (cNSF), which integrates spline layers, as fundamental elements of the invertible function. These concepts are further explained in Sect 3.3.
@@
where $\mathbb{E}(\cdot)$ denotes the expectation operator, and $D_{\text{KL}}(\cdot\;\|\;\cdot)$ represents the Kullback-Leibler divergence. Here we ignore the term $\log p(\bm{\theta}|\mathbf{u}_{o,1:k})$ as it is independent of $\bm{\alpha}$. Following Eq. 11, we write $p_{\bm{\alpha}}(\bm{\theta}|\mathbf{u}_{o,1:k})=p(\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k}))\left|\det\partial\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k})/\partial\bm{\theta}\right|$. Consequently, the above optimization problem can be rewritten as
@@
$\hat{\bm{\bm{\alpha}}}=\operatorname*{arg\,max}_{\bm{\bm{\alpha}}}\frac{1}{M}\sum_{i=1}^{M}\left[\log p(\mathbf{h}_{\bm{\alpha}}(\bm{\theta}^{(i)};\mathbf{u}_{1:k}^{(i)}))+\log|\det\bm{J}_{\mathbf{h}_{\bm{\alpha}}}^{(i)}|\right]$ (15)
@@
$-\log p(\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k}))=\frac{N}{2}\log(2\pi)+\frac{1}{2}\|\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k})\|^{2}$ (17)

### v2

@@ Abstract
We present a likelihood-free probabilistic inversion method based on normalizing flows for high-dimensional inverse problems. The proposed method is composed of two complementary networks: a summary network for data compression and an inference network for parameter estimation. The summary network encodes raw observations into a fixed-size vector of summary features, while the inference network generates samples of the approximate posterior distribution of the model parameters based on these summary features. The posterior samples are produced in a deep generative fashion by sampling from a latent Gaussian distribution and passing these samples through an invertible transformation. We construct this invertible transformation by sequentially alternating conditional invertible neural network and conditional neural spline flow layers. The summary and inference networks are trained simultaneously.
We apply the proposed method to an inversion problem in groundwater hydrology to estimate the posterior distribution of the log-conductivity field conditioned on spatially sparse time-series observations of the system’s hydraulic head responses. The conductivity field is represented with 706 degrees of freedom in the considered problem. The comparison with the likelihood-based iterative ensemble smoother PEST-IES method demonstrates that the proposed method accurately estimates the parameter posterior distribution and the observations’ predictive posterior distribution at a fraction of the inference time of PEST-IES.
@@ 1 Introduction
Inverse problems are key for characterizing parameterized models of physical systems. Unlike forward problems, which predict outcomes given model parameters, inverse problems work in the reverse direction, seeking to estimate model parameters or system states from partial measurements of observable quantities [1]. Inverse problems allow us to update computational models to better align with observed data, thereby enhancing the models’ predictive capabilities [2]. Application areas leveraging inverse problems include medical imaging [3], geophysics [4], structural health monitoring [5], reservoir engineering [6], and astronomy [7], among others.
@@
BI methods can be categorized as follows based on their approach to evaluating the likelihood function that links model parameters to observations: (i) likelihood-based inference and (ii) likelihood-free inference—also known as simulation-based inference. Likelihood-based inference estimates the posterior distribution by explicitly evaluating the likelihood function, which should be explicitly defined and tractable. The gold standard for numerically sampling the posterior distribution is the Markov chain Monte Carlo (MCMC) method. While conceptually straightforward, likelihood-based inference poses various challenges. First, directly evaluating the likelihood function is often either intractable or unfeasible [12] as in the case of multi-level models where sub-models are hierarchically connected [13], continuous-time state-space models [14], and nonlinear dynamic models [15], among others. Second, likelihood-based inference requires the repeated evaluation of the forward model to ensure convergence to the true posterior, which can be computationally intractable depending on the computational cost of the forward model. Third, likelihood-based inference methods are often hampered by the curse of dimensionality, which is potentially challenging for problems with spatially heterogeneous parameters, which often demand high-dimensional characterizations [16]. While the use of surrogate models, such as Gaussian processes [17, 18], polynomial chaos expansion [19, 20], and deep neural network [21, 22] models, can reduce the computational cost, these methods still require likelihood evaluation and introduce additional epistemic uncertainty. Moreover, constructing accurate surrogate models becomes particularly problematic when dealing with high-dimensional inputs and outputs.
@@
In contrast, likelihood-free inference aims to estimate the posterior distribution without directly evaluating the likelihood function and relying solely on synthetic data generated for various parameter choices. Approximate Bayesian computation (ABC) is perhaps the most developed form of likelihood-free inference [23]. In ABC, parameters are repeatedly sampled from a prior distribution, and synthetic data is generated by running the forward model using these sampled parameters. Parameters whose simulated data sufficiently resemble actual observations, as decided by user-defined criteria, are retained as samples of the target posterior; others are discarded. However, ABC has several disadvantages resulting from its accept-reject mechanism: it is challenging to set reasonable criteria for accurate posterior estimation. Additionally, the entire ABC process must be restarted with new measurement datasets, limiting its application to inference for a single set of observed data [24].
@@
In recent developments, deep generative models (DGMs) have emerged as a promising approach for solving inverse problems via likelihood-free inference. Central to DGMs is the strategy of a nonlinear map from a latent distribution, such as a Gaussian, to a complex target distribution. Consequently, approximating the posterior can be done by learning this nonlinear transformation. Such a learning process can be efficiently executed using neural networks trained via back propagation [25]. After sufficient training, one can draw samples from a straightforward latent distribution and produce posterior distribution samples by applying the learned transformation. Numerous studies have explored the use of DGMs for inverse problems, including applications of generative adversarial networks (GANs) [26, 27] and variational autoencoders [28, 29]. GANs and their variants, in particular, have become increasingly popular for likelihood-free inference. Baptista et al. [30] introduced conditional sampling for likelihood-free inference using a monotone GAN, which learns a block triangle transport map between a reference and target measure through adversarial learning combined with a monotonicity penalty. Ramesh et al. [31] developed a conditional GAN to redefine the variational objective for learning an implicit posterior distribution over latent variables, demonstrating significant promise for high-dimensional inverse problems. Additionally, Patel et al. [32] proposed using a GAN to model a field characterized by numerous discrete representations, where the GAN’s generator maps a latent distribution to the target distribution for the field. Despite the significant impact of GANs on posterior estimation, they are notoriously challenging to train, highly sensitive to hyperparameter settings, and prone to the issue of mode collapse [33].
@@
An alternative approach to DGM-based likelihood-free inference is the use of flow-based models, particularly normalizing flows (NFs) [34]. NFs use invertible networks, allowing direct sampling from complex, parameterized probability distributions and precise density evaluation using a change-of-variables formula. NFs have demonstrated strong performance in high-quality posterior estimation across various fields such as economics [35], epidemiology [36], physics [37], computer vision [38], and civil engineering [39]. Additionally, they provide precise log-likelihood evaluations and avoid the training stability issues often encountered in GANs.
@@
In this study, we introduce a likelihood-free inference approach using NFs for estimating high-dimensional spatially heterogeneous parameters in initial-boundary value problems (IBVPs). The proposed model consists of two complementary deep neural networks: a summary network and an inference network. The summary network compresses raw time-series observation data employing one-dimensional convolutions in time, extracting the most informative features, moving beyond manually crafted summary statistics. The inference network is constructed by alternating between two types of flow-based invertible layers: affine coupling layers (ACLs) and spline layers (SLs). By alternating between these two types of layers we enhance the model’s expressivity and flexibility. The summary network forwards the summarized observations to the inference network, thereby conditioning the entire NF on these observations. The inference and summary networks are jointly trained through backpropagation using synthetic data. This concurrent training ensures that the data representation is adequate for estimating the posterior distributions of the parameters. We apply our approach to estimating a hydraulic log-conductivity field in a numerical groundwater model with 706 special computational cells using noisy time-dependent observations of the hydraulic head at 13 sparse locations. The estimation results are compared with the likelihood-based iterative ensemble smoother (IES) method implemented in the PEST software in terms of both accuracy and computational efficiency. The main advantages of the proposed method are as follows:
@@
    It can accurately estimate high-dimensional parameters without the need to evaluate the likelihood function and make any assumptions about the posterior.
@@ 2.1 Inverse problem formulation
The motivating application is estimation of spatially heterogeneous parameters in groundwater flow models. In this work, we consider two-dimensional groundwater flow over the simulation domain $\Omega\in\mathbb{R}^{2}$ governed by the nonlinear partial differential equations
$S_{y}\frac{\partial u(\mathbf{x},t)}{\partial t}=\nabla\cdot[\kappa(\mathbf{x})u(\mathbf{x},t)\nabla u(\mathbf{x},t)]-r(\mathbf{x},t),\quad\mathbf{x}\in\Omega,\,t>0$ (1)
subject to appropriate initial and boundary conditions. Here, $u(\mathbf{x},t)$ denotes the hydraulic head field, $\kappa(\mathbf{x})$ the spatially varying hydraulic conductivity, $r(\mathbf{x},t)$ the source term, and $S_{y}$ the (assumed to be known) specific yield.
For a given $\kappa(\mathbf{x})$, Eq (1) is solved using the control volume finite-difference solver MODFLOW [40]. In the numerical model, the domain is discretized with $N$ cells and the conductivity field is represented with the vector $\mathbf{\kappa}\in\mathbb{R}^{N}$ of its values at the cell centers. We solve the inverse problem for $\mathbf{y}=\ln\bm{\kappa}\in\mathbb{R}^{N}$ rather than $\mathbf{\kappa}$, which guarantees that $\kappa(\mathbf{x})$ is strictly positive.
We assume that only hydraulic head responses $\mathbf{u}$, collected at $N_{u}$ sensor locations ($N_{u}<N$), are available for solving the inverse problem. Because the number of observations is smaller than the number of parameters, this inverse problem is, in general, ill-posed. We note that many inverse methods reduce the dimensionality of the parameter space because of the high computational cost of finding the inverse solution in the full parameter space, an approximation leading to some errors. Here, we formulate the inverse problem in the original $\mathbb{R}^{N}$ space, which makes it possible to solve problems where the dimension reduction is not possible, i.e., when all parameters are uncorrelated and equally important.
The proposed likelihood-free inference model is trained using a synthetic labeled dataset consisting of pairs of hydraulic head values at the sensor locations and $\mathbf{y}$. The $\mathbf{y}$ vectors are generated by sampling the log-conductivity prior distribution, and the corresponding hydraulic head field is computed with MODFLOW.
@@ 2.2 Bayesian inference
Let $\bm{\theta}\coloneqq\mathbf{y}$ denote the uncertain model parameters. Furthermore, let $\mathcal{M}_{k}(\bm{\theta})$ denote the time-series forward model’s output at timestep $k$ consisting of the hydraulic head at $N_{u}$ observation locations. We also consider the excitations, e.g., boundary conditions, initial conditions, and body forces, as fixed in this setting. We model the time-series observations of hydraulic head at sensor locations via the additive noise model
@@
The posterior distribution $p(\bm{\theta}|\mathbf{u}_{o,1:k})$ can be estimated through various Bayesian inference techniques, including MCMC sampling methods and ABC. However, these methods often face challenges such as high computational cost and inefficiency in high-dimensional parameter spaces, as discussed in Sec. 1. We propose a likelihood-free inference approach to address these limitations. Unlike most existing likelihood-free inference approaches, which are typically limited to low-dimensional parameter spaces (e.g., a few or up to a dozen parameter dimensions), our approach leverages the capacity of NFs to infer high-dimensional parameter fields, specifically a 706-dimensional log-conductivity field, using sparse and noisy time-series observations. Our model employs two types of invertible layers—ACLs and SLs—to enhance estimation accuracy. The proposed framework effectively tackles the challenges of estimating high-dimensional, spatially distributed parameter fields in an amortized fashion as described in the following section.
@@ 3 Methodology
The proposed likelihood-free method for parameter estimation consists of three distinct phases, as depicted in Fig. 1. In Phase I, simulation data are generated by sampling model parameters from the prior distribution, $\bm{\theta}\sim p(\bm{\theta})$, and using them to run a forward model. This process produces a training dataset of $M$ pairs, $\{(\bm{\theta}^{(i)},\mathbf{u}^{(i)}_{\text{syn},1:k})\}_{i=1}^{M}$, where $\bm{\theta}^{(i)}$ represents the sampled parameters and $\mathbf{u}^{(i)}_{\text{syn},1:k}$ denotes the corresponding simulated observations over the time steps $1:k$. Phase II consists of offline training the NF model using simulation data. Finally, Phase III consists of online amortized inference, in which we generate samples from the posterior distribution for any observation dataset. The term “online amortized inference” refers to the fact that in our methodology, the intensive computational effort of data simulation and model training are front-loaded during Phases I and II. Consequently, inference can be performed by sampling from the latent simple distribution and the learnt transformation without starting anew for different sets of observation data. This confers a substantial boost in efficiency for applications in which it is necessary to estimate varying parameters (such as boundary conditions) from new observations.
It is worth mentioning that the need for the retraining flow model depends on the nature of the changes to $N_{u}$ (number of observation locations) and $k$ (number of time steps). If the observations contain fewer time steps than the training data, the model does not need to be retrained and can still be used to infer parameters, given that we employ one-dimensional convolutions to summarize the time-series data. If the number of time steps in the observations exceeds those in the training data, the model can also be reused by employing recursive likelihood-free inference [13]. On the other hand, if the number of observation locations differs from that in the training data, retraining the model is necessary. The spatial arrangement of observations directly affects the structure of the likelihood-free inference process and the mappings learnt by the NF model.
Figure 1: The workflow of the proposed method for Bayesian model inversion.
@@ 3.1 Likelihood-free inference by normalizing flow
The proposed likelihood-free inference framework is based on NFs and integrates concepts from Bayesian inference and NF theory. Originally developed by Tabak and Turner [41], and further by Tabak and Vanden-Eijnden [42], NFs were primarily applied to feature extraction and density estimation. NFs are generative models that constructs complex distributions by transforming a known simple probability distribution (e.g., a multivariate normal distribution) through a series of invertible and differentiable transformations.
@@
NF consists of a sequence of $t$ transformations applied to latent variables $\mathbf{z}$ as follows:
@@
$\bm{\theta}=\mathbf{g}_{1}\circ\mathbf{g}_{2}\circ\dots\circ\mathbf{g}_{t}(\mathbf{z}_{0}).$ (7)
Thus, the distribution $p(\bm{\theta})$ can be expressed in terms of the base distribution $p(\mathbf{z}_{0})$ and the Jacobian determinant of the composite transformation $\mathbf{h}(\bm{\theta})$. The complex target distribution is thus approximated by the transformed latent distribution $p(\mathbf{z}_{0})$ using the change-of-variables rule. The distribution $p(\bm{\theta})$ is given by
$p(\bm{\theta})=p(\mathbf{z}_{0})\left|\det\left(\frac{\partial\mathbf{h}(\bm{\theta})}{\partial\bm{\theta}}\right)\right|=p(\mathbf{z}_{0})\left|\det\left(\frac{\partial\mathbf{g}(\mathbf{h}(\bm{\theta}))}{\partial\mathbf{z}_{0}}\right)\right|^{-1}.$ (8)
@@
Figure 2: Schematic of normalizing flow.
@@
$\bm{\theta}\sim p_{\bm{\alpha}}(\bm{\theta}|\mathbf{u}_{o,1:k})\Leftrightarrow\bm{\theta}=\mathbf{h}_{\bm{\alpha}}^{-1}(\mathbf{z};\mathbf{u}_{o,1:k}),\text{ with }\mathbf{z}\sim\mathcal{N}(\mathbf{z}|\bm{0},\mathbf{I}).$ (12)
In practice, the design of invertible functions $\mathbf{g}(\cdot)$ for nonlinear bijective mappings poses significant challenges. Commonly, these functions are parameterized through layers of invertible neural networks, each providing efficient bijections with cost-effective Jacobian evaluations. A detailed review of such architectures is available in [43]. In this work we employ ACLs and SLs as the fundamental elements of the invertible function. These concepts are further explained in Section 3.3.
@@
where $\mathbb{E}(\cdot)$ denotes the expectation operator, and $D_{\text{KL}}(\cdot\;\|\;\cdot)$ represents the KL divergence. Here we ignore the term $\log p(\bm{\theta}|\mathbf{u}_{o,1:k})$ because it is independent of $\bm{\alpha}$. Following Eq. 11, we write $p_{\bm{\alpha}}(\bm{\theta}|\mathbf{u}_{o,1:k})=p(\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k}))\left|\det\partial\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k})/\partial\bm{\theta}\right|$. Consequently, the above optimization problem can be rewritten as
@@
$\hat{\bm{\bm{\alpha}}}=\operatorname*{arg\,max}_{\bm{\bm{\alpha}}}\frac{1}{M}\sum_{i=1}^{M}\left[\log p(\mathbf{h}_{\bm{\alpha}}(\bm{\theta}^{(i)};\mathbf{u}_{1:k}^{(i)}))+\log|\det\bm{J}_{\mathbf{h}_{\bm{\alpha}}}^{(i)}|\right].$ (15)
@@
$-\log p(\mathbf{z}=\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k}))=\frac{N}{2}\log(2\pi)+\frac{1}{2}\|\mathbf{h}_{\bm{\alpha}}(\bm{\theta};\mathbf{u}_{o,1:k})\|^{2}.$ (17)